OpenStudy (anonymous):

Hi How do we solve the following: (using an identity!! the answer is : rad 3 over 2 ) arcsin x - arccos x = pi/6

OpenStudy (anonymous):

Wait. What's the question again? :)

OpenStudy (anonymous):

arcsin x - arccos x = pi/2

OpenStudy (anonymous):

Ooops I meant it equals pi/6

myininaya (myininaya):

\[\arcsin(x)=\arccos(x)+\frac{ \pi}{6}\] \[x=\sin(\arccos(x)+\frac{ \pi}{6})\] \[x=\sin(\arccos(x)) \cos(\frac{\pi}{6})+\sin(\frac{\pi}{6}) \cos(\arccos(x))\] So to write sin(arccos(x)) as an algebraic expression we need to draw a right triangle and use arccos(x)=y => cos(y)=x  so \[\sin(\arccos(x))=\sqrt{1-x^2}\] So we have \[x=\sqrt{1-x^2} \cdot \frac{\sqrt{3}}{2}+\frac{1}{2} \cdot x\] Can you solve this algebraic expression for x? \[x-\frac{1}{2}x=\frac{\sqrt{3}}{2}\sqrt{1-x^2}\] \[\frac{1}{2}{x}=\frac{\sqrt{3}}{2}\sqrt{1-x^2}\] Square both sides \[\frac{1}{4}x^2=\frac{3}{4}(1-x^2)\] \[\frac{1}{4}x^2=\frac{3}{4}-\frac{3}{4}x^2\] I will leave the rest to you :)